OpenStudy (anonymous):

In using induction to prove that 152n + 1 + 32n + 1 is divisible by 18, what statement is represented by S>k? a.) 152^n + 1 + 32^n + 1 = 18p where p is an integer b.)152^n + 1 + 32^n + 1 = p where p is an integer c.)152^n + 1 + 32^n + 1 = 18 d.)152^n + 1 + 32^n + 1/18 e.)18(152^n + 32^n)

OpenStudy (amistre64):

are those exponents by chance? b^(n+1) ??

OpenStudy (amistre64):

other than that, the rest if it looks just as garbled to me. got no idea what S>k would mean in this context

OpenStudy (anonymous):

yes they are! here In using induction to prove that 15^2n + 1 + 3^2n + 1 is divisible by 18, what statement is represented by S>k? a.) 15^2n + 1 + 3^2n + 1 = 18p where p is an integer b.) 15^2n + 1 + 3^2n + 1 = p where p is an integer c.) 15^2n + 1 + 3^2n + 1 = 18 d.) 15^2n + 1 + 3^2n + 1/18 e.) 18(15^2n + 3^2n)

OpenStudy (anonymous):

@amistre64